# Accelerated Parallel Tempering via Neural Transports

### v2

@@ 6 Experiments
In this section, we evaluate our proposed accelerated parallel tempering and its three variants on a variety of targets. We defer all experimental details to Appendix D.
@@ 6.1 Comparison of Acceleration Methods
We evaluate the three variants of APT introduced in Section 5: Flow-APT, CMCD-APT and Diff-APT, against the baseline of PT using the linear path with reference $\pi_{0}={\mathcal{N}}(0,\mathrm{I})$, on a 40-mode Gaussian mixture model in 10 dimensions (GMM-$10$) (Midgley et al., 2023). Table 1 compares the performance of PT vs the APT variants for $N=6,10,30$ chains. We choose these numbers to represent: (1) the smallest $N$ where standard PT obtains round trips (2) the regime $N\approx\Lambda$ where PT is stable, and (3) the regime where $N\gg\Lambda$ and PT is close to optimal. All three APT variants substantially improve over classical PT, with increased round trips with $N$ and $K$. In the small $N$ regime (1) resulted in a 10x to 100x improvement and highlights the potential of APT for challenging problems where PT struggles and parallel chains are constrained. By contrast, when $N\gg\Lambda$ and the improvements over standard PT are less dramatic, they might not offset the cost of additional neural evaluations. Notably, CMCD with $K=5$ and $N=30$ matches the performance of the theoretical limit $T/(2+2\Lambda)\approx 5,349$ for standard PT when $N\to\infty$ as predicted by Theorem 2.
Table 1: PT versus APT with different acceleration methods, targeting a 40-mode Gaussian Mixture model (GMM) target in 10 dimensions and standard Gaussian reference using $N=6,10,30$ parallel chains for $T=100,000$ iterations. For each method, we report the round trips (R), round trips per potential evaluation, denoted as compute-normalised round trips (CN-R), the number of neural network evaluations per parallel chain every iteration, and $\Lambda$ estimated using $N=30$ chains.
@@
Method Neural Call ()($\downarrow$)( ↓ ) $\hat{\Lambda}$ ()($\downarrow$)( ↓ ) R ()($\uparrow$)( ↑ ) CN-R ()($\uparrow$)( ↑ ) R ()($\uparrow$)( ↑ ) CN-R ()($\uparrow$)( ↑ ) R ()($\uparrow$)( ↑ ) CN-R ()($\uparrow$)( ↑ )
@@ 6.2 Scaling with Dimensions
To understand the theoretical performance of $K$-step APT for a fixed number of chains when we scale the dimension $d$, we use the fact that the path of distributions $(\pi^{\text{VP}}_{s})_{s\in[0,1]}$ induced by a VP diffusion process initialised at GMM-$d$ is analytically tractable, in order to run $K$-step Diff-APT with the true diffusion path. In Figure 2, we compare the resulting round trip rate and compute-normalised round trip rate of the different algorithms when we fix the number of chains to 30.
We see that, for all values of $d$, the round trip rate of Diff-APT monotonically increases over PT as we increase $K$, with greater gains for larger values of $d$, demonstrating that the accelerated swap improves the communication of states over the standard PT swap. We see a substantial improvement in round trip rate (right) with acceleration compared to PT ($K=0$) and a minor difference between algorithms as $K$ increases when normalised for compute (Right), suggesting the extra computation for Diff-APT is justified.
@@ 6.3 Free Energy Estimator
As we described in Section 3.2, a by-product of Algorithm 1 is the estimation of change in free energy $\Delta F$. We compare free energy estimates from CMCD-APT and Diff-APT against PT on two targets: DoubleWell(DW)-4, a particle system in Cartesian coordinates; and ManyWell(MW)-32, a highly multi-modal density introduced by Midgley et al. (2023). Figure 3 presents box-plots of $30$ free energy estimates for DW-4 and MW-32 using 1,000 samples each.
Several key observations emerge: (1) Both CMCD-APT and Diff-APT exhibit markedly lower variance and bias than classical PT, across both targets. (2) For CMCD/Diff-APT, the variance and bias reduce steadily as $K$ increases. (3) For both PT and CMCD/Diff-APT, the variance and bias markedly decrease as the number of chains $N$ increases, thereby empirically validating Proposition 1.
@@ 6.4 Comparing APT with Neural Samplers
A significant advantage of APT is its asymptotic consistency: unlike most neural samplers, the Metropolis correction ensures APT does not incur a bias if the neural proposal is not poorly trained. In Figure 4, we demonstrate this property for both Diff-APT and CMCD-APT. Specifically, for CMCD-APT and Diff-APT, we take the trained model and map samples directly from the reference distribution to the target by concatenating the learned transports between each successive pair of intermediate distributions, labelled as CMCD and Diffusion respectively in Figure 4.
@@ A.1 Further Details of Round Trips
Consider the APT Markov chain $\mathbf{X}_{t}=(X^{0}_{t},\ldots,X^{N}_{t})$ constructed by Algorithm 1. Assume there are $m=0,\ldots,N$ parallel threads, each storing a component of $\mathbf{X}_{t}$ at time $t$. During the communication step, it is equivalent to swap indices rather than states. This has computational advantages in a distributed implementation of PT since swapping states across machines incurs a cost of $O(d)$ compared to the $O(1)$ cost of swapping indices. By tracking how the indices are shuffled, we can recover the PT state and also track the communication between reference and target through the dynamics of the permutated indices. See [Syed et al., 2022, Surjanovic et al., 2023, Algorithm 5] for details on implementing PT with a distributed implementation.
Summarising Syed et al. [2022], we define the index process $(\mathbf{I}_{t},\bm{\epsilon}_{t})$ where $\mathbf{I}_{t}=(I^{0}_{t},\ldots,I^{N}_{t})$ is a sequence of permutations of $[N]=\{0,\ldots,N\}$ that tracks the underlying communication of states in $\mathbf{X}_{t}$, and $\bm{\epsilon}_{t}=(\epsilon^{0}_{t},\ldots,\epsilon^{N}_{t})$ with $\epsilon^{m}_{t}\in\{-1,1\}$ tracks the direction of the swap proposal on machine $m$.
We initialize $I^{m}_{0}=m$ and $\epsilon^{m}_{0}=-1$ if $m$ is even and $\epsilon^{m}_{0}=1$ if $m$ is odd. The subsequent values of $\mathbf{I}_{t}$ are then determined by the swap moves. At iteration $t$ of PT, we apply the same swaps to the components of $\mathbf{I}_{t}$ that are proposed and accepted during the communication phase. As discussed in Syed et al. [2022], for non-reversible communication, $I^{m}_{t}$ and $\epsilon^{m}_{t}$ satisfy the recursion:
$I^{m}_{t}$ =$=\begin{cases}I^{m}_{t-1}+\epsilon^{m}_{t-1},&\text{if }S^{I^{m}_{t-1}\vee I^{m}_{t-1}+\epsilon^{m}_{t-1}}_{t}=1\\ I^{m}_{t-1},&\text{if }S^{I^{m}_{t-1}\vee I^{m}_{t-1}+\epsilon^{m}_{t-1}}_{t}=0\end{cases},$
$\epsilon^{m}_{t}$ =$=\begin{cases}\epsilon^{m}_{t-1},&\text{if }I^{m}_{t}=I^{m}_{t-1}+\epsilon^{m}_{t-1}\\ -\epsilon^{m}_{t-1},&\text{if }I^{m}_{t}=I^{m}_{t-1}\end{cases}.$
@@
Hence, for a realization of $\mathbf{X}_{t}$ with $T$ steps, we define the round trips for index $n\in[N]$ as the number of times the index $n$ in $\mathbf{I}_{t}$ completes the journey from the 00-th component to the $N$-th component and then back to the 00-th component—i.e., completes the round trip from the reference to the target and back again.
@@
Notably, $T^{m}_{\downarrow,j}$ represents the $j$-th time the index for machine $m$ has traversed from 00 to $N$ and back to 00, thus defining a round trip. A round trip indicates that a sample on machine $m$ from the reference propagated a reference sample to the target independent of previously visited target states on the same machine. Let $R^{m}_{T}=\max\{j:T^{m}_{\downarrow,j}\leq T\}$ denote the number of round trips that occur on machine $m$ by iteration $T$. The overall round trips count is then defined as the sum of round trips over all index values: $R_{T}=\sum_{m=0}^{N}R^{m}_{T}$. Finally, the empirical round trip rate $\tau_{T}=R_{T}/T$ represents the fraction of PT iterations during which a round trip occurred, and our objective is to maximise the expected round trip rate $\tau$ as $T\to\infty$:
@@
A.2 Proof of Theorem 1
A.2.1 Proof of Ergodicity (Theorem 1, Part 1)
@@
At stationary, we have $(X_{t}^{n-1},X_{t}^{n})\sim\pi^{n-1}\otimes\pi^{n}$. We next simulate the two paths $\accentset{\rightharpoonup}{X}_{t,0:K}^{n-1}=(\accentset{\rightharpoonup}{X}^{n-1}_{t,0},\dots,\accentset{\rightharpoonup}{X}^{n-1}_{t,K})$ and $\accentset{\leftharpoonup}{X}_{t,0:K}^{n}=(\accentset{\leftharpoonup}{X}^{n}_{t,0},\dots,\accentset{\leftharpoonup}{X}^{n}_{t,K})$ as described in Section 3 and calculate the acceptance probability
@@
given in Section 3.2. The new states $\hat{X}_{t}^{n-1},\hat{X}_{t}^{n}$ are determined by
$(\hat{X}_{t}^{n-1},\hat{X}_{t}^{n})=\begin{cases}(\accentset{\leftharpoonup}{X}_{t,0}^{n},\accentset{\rightharpoonup}{X}_{t,K}^{n-1})&\textrm{ with probability }\alpha,\\ (\accentset{\rightharpoonup}{X}_{t,0}^{n-1},\accentset{\leftharpoonup}{X}_{t,K}^{n})&\textrm{ with probability }1-\alpha.\end{cases}$
@@
and to apply an instance of MCMC algorithms with a deterministic proposal [Tierney, 1998], here given by $\mathfrak{T}$. For completeness we give a self-contained proof. Write
@@
To arrive at Equation 8 we need to show that
@@
Thus Equation 9 is established.
Finally, the accelerated-PT Markov chain is aperiodic because the swaps can be rejected. The chain is clearly irreducible if each exploration kernel is irreducible. (In fact it can be proved, using more complicated arguments, that the chain is irreducible if the exploration kernel for the reference is irreducible and, for all $n$, the two distributions ${\mathbb{P}}_{K}^{n-1}$ and ${\mathbb{Q}}_{K}^{n}$ are mutually absolutely continuous.) By Roberts and Rosenthal [2004, Theorem 4, Fact 5], the Markov chain is ergodic and in particular the law of large numbers hold. ∎
A.2.2 Proof of Rejection Rate (Theorem 1, Part 2)
@@
From Equation 10, Equation 11, and Lemma 1 below we have
@@
Combining Equation 12 and Equation 13 we get the desired result. ∎
A.2.3 Proof of Proposition 1
@@
We will show the consistency of the expectation and free energy estimator separately. By Theorem 1, the PT chain $\textbf{X}_{t}=(X^{0}_{t},\ldots,X^{N}_{t})$ is ergodic with stationary distribution $\pi^{0}\otimes\cdots\otimes\pi^{N}$. By the ergodic theorem, we have
@@
We will now show the consistency of the free energy estimator. By definition of $W^{n}_{K}$ in Equation 3, we can write $\Delta F_{n}$ in terms of expectations of $W^{n}_{K}$ and the Radon-Nikodym derivative between ${\mathbb{P}}^{n-1}_{K}$ and ${\mathbb{Q}}^{n}_{K}$ respectively:
@@ A.3 Accelerated PT as Vanilla PT on Extended Space
which is straightforward from Equation 15. ∎
@@
  • In addition, accelerated PT inserts $N-1$ distributions between $\pi_{\operatorname{ex}}^{0}$ and $\pi_{\operatorname{ex}}^{N}$.

### v3

@@ 6 Experiments
In this section, we evaluate our proposed accelerated parallel tempering and its three variants on a variety of targets. We defer all experimental details to Appendix˜D.
@@ 6.1 Comparison of Acceleration Methods
We evaluate the three variants of APT introduced in Section˜5: Flow-APT, CMCD-APT and Diff-APT, against the baseline of PT using the linear path with reference $\pi_{0}={\mathcal{N}}(0,\mathrm{I})$, on a 40-mode Gaussian mixture model in 10 dimensions (GMM-$10$) (Midgley et al., 2023). Table˜1 compares the performance of PT vs the APT variants for $N=6,10,30$ chains. We choose these numbers to represent: (1) the smallest $N$ where standard PT obtains round trips (2) the regime $N\approx\Lambda$ where PT is stable, and (3) the regime where $N\gg\Lambda$ and PT is close to optimal. All three APT variants substantially improve over classical PT, with increased round trips with $N$ and $K$. In the small $N$ regime (1) resulted in a 10x to 100x improvement and highlights the potential of APT for challenging problems where PT struggles and parallel chains are constrained. By contrast, when $N\gg\Lambda$ and the improvements over standard PT are less dramatic, they might not offset the cost of additional neural evaluations. Notably, CMCD with $K=5$ and $N=30$ matches the performance of the theoretical limit $T/(2+2\Lambda)\approx 5,349$ for standard PT when $N\to\infty$ as predicted by Theorem˜2.
[…]
@@
Method Neural Call $(\downarrow)$ $\hat{\Lambda}$ $(\downarrow)$ R $(\uparrow)$ CN-R $(\uparrow)$ R $(\uparrow)$ CN-R $(\uparrow)$ R $(\uparrow)$ CN-R $(\uparrow)$
@@ 6.2 Scaling with Dimensions
To understand the theoretical performance of $K$-step APT for a fixed number of chains when we scale the dimension $d$, we use the fact that the path of distributions $(\pi^{\text{VP}}_{s})_{s\in[0,1]}$ induced by a VP diffusion process initialised at GMM-$d$ is analytically tractable, in order to run $K$-step Diff-APT with the true diffusion path. In Figure˜2, we compare the resulting round trip rate and compute-normalised round trip rate of the different algorithms when we fix the number of chains to 30. For all values of $d$, the round trip rate of Diff-APT monotonically increases over PT as we increase $K$, with greater gains for larger values of $d$, demonstrating that the accelerated swap improves the communication of states over the standard PT swap. We see a substantial improvement in round trip rate (right) with acceleration compared to PT ($K=0$) and a minor difference between algorithms as $K$ increases when normalised for compute (Right), suggesting the extra computation for APT is justified.
@@ 6.3 Free Energy Estimator
As we described in Section˜3.2, a by-product of Algorithm˜1 is the estimation of change in free energy $\Delta F$. We compare free energy estimates from CMCD-APT and Diff-APT against PT on two targets: DoubleWell(DW)-4, a particle system in Cartesian coordinates; and ManyWell(MW)-32, a highly multi-modal density introduced by Midgley et al. (2023). Figure˜3 presents box-plots of $30$ free energy estimates for DW-4 and MW-32 using 1,000 samples each.
Several key observations emerge: (1) Both CMCD-APT and Diff-APT exhibit markedly lower variance and bias than classical PT, across both targets. (2) For CMCD/Diff-APT, the variance and bias reduce steadily as $K$ increases. (3) For both PT and CMCD/Diff-APT, the variance and bias markedly decrease as the number of chains $N$ increases, thereby empirically validating Proposition˜1.
@@ 6.4 Comparing APT with Neural Samplers
A significant advantage of APT is its asymptotic consistency: unlike most neural samplers, the Metropolis correction ensures APT does not incur a bias if the neural proposal is poorly trained. In Figure˜4, we demonstrate this property for both Diff-APT and CMCD-APT. Specifically, for CMCD-APT and Diff-APT, we take the trained model and map samples directly from the reference distribution to the target by concatenating the learned transports between each successive pair of intermediate distributions, labelled as CMCD and Diffusion respectively in Figure˜4.
@@
Table 2: PT versus CMCD-APT targeting Alanine Dipeptide. We use the same metrics as defined in Table 1.
Method $\hat{\Lambda}$ $(\downarrow)$ R $(\uparrow)$ CN-R $(\uparrow)$
CMCD-APT ($K=1$) 3.23 465 232.5
CMCD-APT ($K=2$) 3.15 597 199
CMCD-APT ($K=5$) 3.09 627 104.5
PT 3.38 199 99.5
6.5 Alanine Dipeptide
Finally, to provide a realistic target on which to evaluate APT, we take the Boltzmann distribution of Alanine Dipeptide in Cartesian coordinates at 300K, where we compare PT against CMCD-APT with $N=4$. This is a small molecule with 22 atoms (66 dimensions in total) which is highly challenging to sample from, due to many prohibitive energy barriers and high multimodality. Following the convention of molecular dynamics, we take the reference distribution to be a tempered version of the target, corresponding to 1200K. We run each algorithm for $T=50,000$ iterations and report the results in Table 2. We can see that CMCD-APT is able to provide significant acceleration in this realistic setting.
@@ A.1 Further Details of Round Trips
Consider the APT Markov chain $\mathbf{X}_{t}=(X^{0}_{t},\ldots,X^{N}_{t})$ constructed by Algorithm˜1. Assume there are $m=0,\ldots,N$ parallel threads, each storing a component of $\mathbf{X}_{t}$ at time $t$. During the communication step, it is equivalent to swap indices rather than states. This has computational advantages in a distributed implementation of PT since swapping states across machines incurs a cost of $O(d)$ compared to the $O(1)$ cost of swapping indices. By tracking how the indices are shuffled, we can recover the PT state and also track the communication between reference and target through the dynamics of the permutated indices. See (Syed et al., 2022; Surjanovic et al., 2023, Algorithm 5) for details on implementing PT with a distributed implementation.
Summarising Syed et al. (2022), we define the index process $(\mathbf{I}_{t},\bm{\epsilon}_{t})$ where $\mathbf{I}_{t}=(I^{0}_{t},\ldots,I^{N}_{t})$ is a sequence of permutations of $N]=\{0,\ldots,N\}$ that tracks the underlying communication of states in $\mathbf{X}_{t}$, and $\bm{\epsilon}_{t}=(\epsilon^{0}_{t},\ldots,\epsilon^{N}_{t})$ with $\epsilon^{m}_{t}\in\{-1,1\}$ tracks the direction of the swap proposal on machine $m$.
We initialize $I^{m}_{0}=m$ and $\epsilon^{m}_{0}=-1$ if $m$ is even and $\epsilon^{m}_{0}=1$ if $m$ is odd. The subsequent values of $\mathbf{I}_{t}$ are then determined by the swap moves. At iteration $t$ of PT, we apply the same swaps to the components of $\mathbf{I}_{t}$ that are proposed and accepted during the communication phase. As discussed in Syed et al. (2022), for non-reversible communication, $I^{m}_{t}$ and $\epsilon^{m}_{t}$ satisfy the recursion:
$I^{m}_{t}$ $=\begin{cases}I^{m}_{t-1}+\epsilon^{m}_{t-1},&\text{if }S^{I^{m}_{t-1}\vee I^{m}_{t-1}+\epsilon^{m}_{t-1}}_{t}=1\\ I^{m}_{t-1},&\text{if }S^{I^{m}_{t-1}\vee I^{m}_{t-1}+\epsilon^{m}_{t-1}}_{t}=0\end{cases},$
$\epsilon^{m}_{t}$ $=\begin{cases}\epsilon^{m}_{t-1},&\text{if }I^{m}_{t}=I^{m}_{t-1}+\epsilon^{m}_{t-1}\\ -\epsilon^{m}_{t-1},&\text{if }I^{m}_{t}=I^{m}_{t-1}\end{cases}.$
@@
Hence, for a realization of $\mathbf{X}_{t}$ with $T$ steps, we define the round trips for index $n\in[N]$ as the number of times the index $n$ in $\mathbf{I}_{t}$ completes the journey from the 0-th component to the $N$-th component and then back to the 0-th component—i.e., completes the round trip from the reference to the target and back again.
@@
Notably, $T^{m}_{\downarrow,j}$ represents the $j$-th time the index for machine $m$ has traversed from 0 to $N$ and back to 0, thus defining a round trip. A round trip indicates that a sample on machine $m$ from the reference propagated a reference sample to the target independent of previously visited target states on the same machine. Let $R^{m}_{T}=\max\{j:T^{m}_{\downarrow,j}\leq T\}$ denote the number of round trips that occur on machine $m$ by iteration $T$. The overall round trips count is then defined as the sum of round trips over all index values: $R_{T}=\sum_{m=0}^{N}R^{m}_{T}$. Finally, the empirical round trip rate $\tau_{T}=R_{T}/T$ represents the fraction of PT iterations during which a round trip occurred, and our objective is to maximise the expected round trip rate $\tau$ as $T\to\infty$:
@@
A.2 Proof of Theorem˜1
A.2.1 Proof of Ergodicity (Theorem˜1, Part 1)
@@
At stationary, we have $(X_{t}^{n-1},X_{t}^{n})\sim\pi^{n-1}\otimes\pi^{n}$. We next simulate the two paths $\accentset{\rightharpoonup}{X}_{t,0:K}^{n-1}=(\accentset{\rightharpoonup}{X}^{n-1}_{t,0},\dots,\accentset{\rightharpoonup}{X}^{n-1}_{t,K})$ and $\accentset{\leftharpoonup}{X}_{t,0:K}^{n}=(\accentset{\leftharpoonup}{X}^{n}_{t,0},\dots,\accentset{\leftharpoonup}{X}^{n}_{t,K})$ as described in Section˜3 and calculate the acceptance probability
@@
given in Section˜3.2. The new states $\hat{X}_{t}^{n-1},\hat{X}_{t}^{n}$ are determined by
[…]
@@
and to apply an instance of MCMC algorithms with a deterministic proposal (Tierney, 1998), here given by $\mathfrak{T}$. For completeness we give a self-contained proof. Write
@@
To arrive at Equation˜8 we need to show that
@@
Thus Equation˜9 is established.
Finally, the accelerated-PT Markov chain is aperiodic because the swaps can be rejected. The chain is clearly irreducible if each exploration kernel is irreducible. (In fact it can be proved, using more complicated arguments, that the chain is irreducible if the exploration kernel for the reference is irreducible and, for all $n$, the two distributions ${\mathbb{P}}_{K}^{n-1}$ and ${\mathbb{Q}}_{K}^{n}$ are mutually absolutely continuous.) By Roberts & Rosenthal (2004, Theorem 4, Fact 5), the Markov chain is ergodic and in particular the law of large numbers hold. ∎
A.2.2 Proof of Rejection Rate (Theorem˜1, Part 2)
@@
From Equation˜10, Equation˜11, and Lemma˜1 below we have
@@
Combining Equation˜12 and Equation˜13 we get the desired result. ∎
A.2.3 Proof of Proposition˜1
@@
We will show the consistency of the expectation and free energy estimator separately. By Theorem˜1, the PT chain $\textbf{X}_{t}=(X^{0}_{t},\ldots,X^{N}_{t})$ is ergodic with stationary distribution $\pi^{0}\otimes\cdots\otimes\pi^{N}$. By the ergodic theorem, we have
@@
We will now show the consistency of the free energy estimator. By definition of $W^{n}_{K}$ in Equation˜3, we can write $\Delta F_{n}$ in terms of expectations of $W^{n}_{K}$ and the Radon-Nikodym derivative between ${\mathbb{P}}^{n-1}_{K}$ and ${\mathbb{Q}}^{n}_{K}$ respectively:
@@ A.3 Accelerated PT as Vanilla PT on Extended Space
which is straightforward from Equation˜15. ∎
@@
  • In addition, accelerated PT inserts N1N-1 distributions between $\pi_{\operatorname{ex}}^{0}$ and $\pi_{\operatorname{ex}}^{N}$.